i dont understand how to solve quadratic equations by completing the square root.
im in algebra ags chapter 13
problem: ((( 2s^2+5s=3 )))

Answer by stanbon(75887)

i dont understand how to solve quadratic equations by completing the square root.
im in algebra ags chapter 13
problem: 2s^2+5s = 3
-----
Factor out the "2" on the left side:
---
2(s^2+(5/2)s+?) = 3 + 2*?
-----
Complete the square on the s-terms.
-------
2(s^2+(5/2)s+(5/4)^2) = 3 + 2*(5/4)^2
--------
Simplify:
2(s+(5/4))^2 = 3+(50/16) = 98/16
------------------
Divide both sides by 2 to get:
(s+(5/4))^2 = 49/16
----
Take the square root of both sides:
s+(5/4) = +7/4 or s+(5/4)= -(7/4)
---
Solve for "s":
s = 1/2 or s = -3
=========================
Cheers,
Stan H.